# Superconductivity and Low Energy Excitations in an Attractive Hubbard Model

### v1

@@ 10 Spectral Gap of Single-Fermion Excitations
$\frac{\langle\Phi_{\rm GS},\tilde{A}_{\uparrow}^{\dagger}\tilde{A}_{\uparrow}\Phi_{\rm GS}\rangle}{\|\Phi_{\rm GS}\|^{2}}\leq\frac{\langle\Phi_{\rm GS},\tilde{A}_{\uparrow}^{\dagger}\tilde{A}_{\uparrow}\Phi_{\rm GS}\rangle}{\|\Phi_{\rm GS}\|^{2}}+\frac{\langle\Phi_{\rm GS},\tilde{A}_{\uparrow}\tilde{A}_{\uparrow}^{\dagger}\Phi_{\rm GS}\rangle}{\|\Phi_{\rm GS}\|^{2}}=\sum_{x\in\Lambda}|\tilde{\alpha}_{x}|^{2}.$ (10.22)
@@
Following Lieb’s theorem [4], for any even number $N>0$ of the fermions, the ground state is unique and shows spin singlet. Therefore, the number $N_{\uparrow}$ of fermions with up spin satisfies $N_{\uparrow}=N_{\downarrow}$ and $N=N_{\uparrow}+N_{\downarrow}$. Clearly, the filling condition $\nu_{\downarrow}<1/8$ can be realized.
Our result is summarized as follows:
Theorem 10.1
Let $\Phi_{\rm GS}$ be the ground state of the translationally invariant Hamiltonian $H^{(\Lambda)}$ of (10.1) on the lattice $\Lambda\subset\mathbb{Z}^{d}$ with an even number $N>0$ of fermions. Assume that the filling factor $\nu_{\downarrow}:=N_{\downarrow}/|\Lambda|$ for the number $N_{\downarrow}$ of fermions with down spin satisfies $\nu_{\downarrow}<1/8$. Then, the single-fermion excitation $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ with the operator $\mathcal{A}_{\uparrow}$ of (10.2) has a non-vanishing spectral gap above the ground state $\Phi_{\rm GS}$ for a sufficiently small $|t|/U$ if the excitation $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ satisfies the norm condition (10.5).
11 The Case of Lieb Lattice
@@
Let $\Phi_{\rm GS}$ be the ground state of the translationally invariant Hamiltonian $H^{(\Lambda)}$ of (10.1) on the Lieb lattice $\Lambda$ with an even number $N>0$ of fermions. The Lieb lattice has the above site symmetry and satisfies the condition (11.1) with $0<a<1/2$. Assume that the filling factor $\nu_{\downarrow}:=N_{\downarrow}/|\Lambda|$ for the number $N_{\downarrow}$ of fermions with down spin satisfies $\nu_{\downarrow}<a/8$. Then, the single-fermion excitation $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ with the operator $\mathcal{A}_{\uparrow}$ of (10.2) has a non-vanishing spectral gap above the ground state $\Phi_{\rm GS}$ for a sufficiently small $|t|/U$ if the excitation $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ satisfies the norm condition (10.5).
@@ 12 Low-Lying Excitations
In this section, we require lattice translational invariance for the Hamiltonian $H^{(\Lambda)}$. We introduce
@@
where $p$ is a wavenumber vector. In order to evaluate the energy of low-lying excitations above a ground state $\Phi_{\rm GS}$, we consider a trial state $O_{\eta}(p)\Phi_{\rm GS}$. The variational enegy is given by
@@
Assume lattice translational invariance for the Hamiltonian $H^{(\Lambda)}$ of the attractive Hubbard model. If the number of fermions in a ground state $\Phi_{\rm GS}$ deviates from that of the half filling by the order of the volume, then the following bound is vaild:
@@
Since we can choose the ground state $\Phi_{\rm GS}$ such that $\Lambda-N$ is positive and of order $|\Lambda|$. By combining this with the above bound (12.11), we obtain the desired bound (12.3) for the excitation energy $\Delta E$ of (12.2).
@@
Assume lattice translational invariance and the above spatial inversion symmetry for the Hamiltonian $H^{(\Lambda)}$ of the attractive Hubbard model. If the number of fermions in a ground state $\Phi_{\rm GS}$ deviates from that of the half filling by the order of the volume, then the following bound is vaild:

### v2

@@ 6.1 The Case of Hypercubic Lattice
$\frac{\langle\Phi_{\rm GS},\tilde{\mathcal{A}}_{\uparrow}^{\dagger}\tilde{\mathcal{A}}_{\uparrow}\Phi_{\rm GS}\rangle}{\|\Phi_{\rm GS}\|^{2}}\leq\frac{\langle\Phi_{\rm GS},\tilde{\mathcal{A}}_{\uparrow}^{\dagger}\tilde{\mathcal{A}}_{\uparrow}\Phi_{\rm GS}\rangle}{\|\Phi_{\rm GS}\|^{2}}+\frac{\langle\Phi_{\rm GS},\tilde{\mathcal{A}}_{\uparrow}\tilde{\mathcal{A}}_{\uparrow}^{\dagger}\Phi_{\rm GS}\rangle}{\|\Phi_{\rm GS}\|^{2}}=\sum_{x\in\Lambda}|\tilde{\alpha}_{x}|^{2}.$ (6.22)
@@
6.2 The Case of Lieb Lattice
@@
Let $\Phi_{\rm GS}$ be the ground state of the translationally invariant Hamiltonian $H^{(\Lambda)}$ of (6.1) on the Lieb lattice $\Lambda$ with an even number $N>0$ of fermions, and let $\mathcal{A}_{\uparrow}$ be given by (6.2). Here $\Lambda$ has the above site symmetry and satisfies (6.28) with $0<a<1/2$. Assume that the filling factor $\nu_{\downarrow}:=N_{\downarrow}/|\Lambda|$ satisfies $\nu_{\downarrow}<a/8$. Then, for a sufficiently small $|t|/U$, the excitation energy $\Delta(\mathcal{A}_{\uparrow}\Phi_{\rm GS})$ is bounded from below by a positive constant, if the state $\mathcal{A}_{\uparrow}\Phi_{\rm GS}$ satisfies the norm condition (6.5).
@@ 7 Low-Lying Excitations
In this section, we discuss gapless excitations above the ground state. Our aim is to obtain an upper bound on the excitation energy $\Delta E(p)$ for momenta $p$ under certain assumptions. For Heisenberg models, spin-wave excitations have been widely studied, whereas analogous results for the Hubbard model seem to be limited. In the Hubbard model, Momoi [10], following the method of [14, 18, 1], obtained related bounds for the Nambu–Goldstone modes in the magnetic and superconducting phases when the symmetries are spontaneously broken, together with additional assumptions on the dynamical structure factor. In the present paper, we work on finite systems and construct a spin-wave excitation using the deviation of the number of fermions from half filling, without assuming either a non-vanishing order parameter or any conditions on structure factors. We emphasize that our estimates rely essentially only on commutation relations, and do not require any further assumptions beyond the filling condition stated below. It is also worth noting that establishing orthogonality of low energy excitations above infinite-volume ground states calls for careful treatment; see [4].
The argument is independent of the topics in the previous sections, and we impose lattice translational invariance on the Hamiltonian $H^{(\Lambda)}$. We introduce
@@
where $p$ is a wavenumber vector. In order to evaluate the energy of low-lying excitations above a ground state $\Phi_{\rm GS}$, we consider a trial state $O_{\eta}(p)\Phi_{\rm GS}$. The variational energy is given by
@@
Assume lattice translational invariance for the Hamiltonian $H^{(\Lambda)}$ of the attractive Hubbard model. If the number of fermions, $N$, in a ground state $\Phi_{\rm GS}$ satisfies that $|N-|\Lambda||\geq\delta_{0}|\Lambda|$ for some $\delta_{0}>0$, then the following bound is valid:
@@
Since we can choose the ground state $\Phi_{\rm GS}$ such that $|\Lambda|-N$ is positive and of order $|\Lambda|$, by combining this with the above bound (7.11), we obtain the desired bound (7.3) for the excitation energy $\Delta E$ of (7.2).
@@
Assume lattice translational invariance and the above spatial inversion symmetry for the Hamiltonian $H^{(\Lambda)}$ of the attractive Hubbard model. If the number of fermions, $N$, in a ground state $\Phi_{\rm GS}$ satisfies that $|N-|\Lambda||\geq\delta_{0}|\Lambda|$ for some $\delta_{0}>0$, then the following bound is valid: